# Attention Overlap Is Responsible for The Entity Missing Problem in Text-to-image Diffusion Models!

### v1

@@ Appendix B Compositional Evaluation Metrics
VQA Score: Through the visual question answering (VQA) evaluation method, for each entity in a given pair of prompts and generated images, we prepare a question, “Is there any {entity} in the image?” and give it to the VQA model along with the corresponding image. If the VQA model answers “yes”, we consider a score of $1$; otherwise, it is 00. Similar to human evaluation, we proposed two options of soft and harsh averaging in such a way that in a harsh scenario, the overall score of a prompt is $1$, if and only if the VQA answers to all questions are “yes”. At the end, we average all the prompt scores. Notably, BLIP-VQA Li et al. [2022] and TIFA Hu et al. [2023] models are used for VQA evaluation.
Captioning Score: In caption evaluation, we use an image captioning model called BLIP-Captioner [Li et al., 2022] to produce a caption for each generated image. Then, the cosine similarity between the caption and the corresponding original prompt in CLIP [Radford et al., 2021] embedding space is considered as the score of that prompt, and the final score is calculated by averaging across prompt scores.
CLIP Similarity: In this evaluation method, we calculate the cosine similarity between a prompt and its corresponding generated image in CLIP’s shared embedding space, treating it as the score [Hessel et al., 2021]. Similar to other metrics, Averaging these scores across prompts yields the final score.
@@ C.1 State-of-the-art Loss Functions
The proposed loss functions of three well-knwon training-free approach, Attant-and-Excite [Chefer et al., 2023], Divide-and-Bind [Li et al., 2023], and Predicated-Diffusion [Sueyoshi and Matsubara, 2024], are demonstrated in equations 8, 9, and 10, respectively. At each denoising step $t$, the latent embedding $z_{t}$ is modified in the direction opposite to the gradient of these loss functions with respect to $z_{t}$.
@@ D.2 Results of Free-Style Prompts
In the previous experiments, prompts followed a fixed format of “a {entity} and a {entity}”. To evaluate the generalizability of our method to more diverse linguistic structures, we also tested it on free-form captions from the COCO dataset [Lin et al., 2014]. Table 10 presents the results for COCO captions containing at least three entities, utilizing Stable Diffusion 1.4 as the backbone model. The overlap-based loss functions demonstrate significant improvements over the baseline methods.
@@ D.3 Results of Other Backbones
In addition to SD v1.4, we assessed our approach on other recent backbones, SD v2-base, SD v2 [Rombach et al., 2022], and SD-XL [Podell et al., 2023], to validate its effectiveness across different model sizes. Tables 11 and 12 present the results on two-entity and three-entity prompts from our COCO-Comp dataset using Stable Diffusion v2-base. Moreover, tables 13 and 14 demonstrate the obtained results on two-entity prompts of the COCO-Comp dataset, using Stable Diffusion v2 and Stable Diffusion XL, respectively. Across all the backbones, the overlap-based methods outperform other baselines in all evaluation metrics. This suggests that our overlap-based approach can generalize well to different diffusion model architectures and sizes.
@@ Appendix E Additional Qualitative Results
Figures 7, 8, 9, and 10 present additional qualitative comparisons of images generated between our overlap-based methods and various baselines on two-object prompts from our COCO-Comp dataset using Stable Diffusion XL, Stable Diffusion v2, Stable Diffusion v1.4, Stable Diffusion v2-base, respectively.
Figure 7: Qualitative comparison between the proposed overlap-based methods and the state-of-the-art baselines on diverse prompts from our COCO-Comp dataset using Stable Diffusion XL.
@@ Appendix G Limitations
The poor capability of the CLIP text encoder [Radford et al., 2021] in understanding and representing the compositionality of complex texts [Yuksekgonul et al., 2023, Thrush et al., 2022] may have a significant impact on the performance of T2I models. Since most of these models rely on the CLIP text encoder as their backbone for processing textual inputs, any limitations or weaknesses in the encoder’s ability to capture and encode the compositional structure and semantics of the text can propagate through the entire system, leading to suboptimal results or failure modes in the generated images. Surprisingly, most of the related works have not paid sufficient attention to this potential root cause of the failures. Instead, they often focus on proposing solutions without thoroughly investigating the underlying reasons behind the problems. They tend to attribute the issues to other factors without considering the limitations of the CLIP text encoder itself. To address this, we plan to focus our future work on investigating and improving the text encoding capabilities of T2I models, specifically targeting their ability to handle compositional complexities in text prompts.

### v2

@@ Appendix B Compositional Evaluation Metrics
VQA Score: Through the visual question answering (VQA) evaluation method, for each entity in a given pair of prompts and generated images, we prepare a question, “Is there any {entity} in the image?” and give it to the VQA model along with the corresponding image. If the VQA model answers “yes”, we consider a score of $1$; otherwise, it is 00. Similar to human evaluation, we proposed two options of soft and harsh averaging in such a way that in a harsh scenario, the overall score of a prompt is $1$, if and only if the VQA answers to all questions are “yes”. At the end, we average all the prompt scores. Notably, BLIP-VQA Li et al. (2022) and TIFA Hu et al. (2023) models are used for VQA evaluation.
Captioning Score: In caption evaluation, we use an image captioning model called BLIP-Captioner (Li et al., 2022) to produce a caption for each generated image. Then, the cosine similarity between the caption and the corresponding original prompt in CLIP (Radford et al., 2021) embedding space is considered as the score of that prompt, and the final score is calculated by averaging across prompt scores.
CLIP Similarity: In this evaluation method, we calculate the cosine similarity between a prompt and its corresponding generated image in CLIP’s shared embedding space, treating it as the score (Hessel et al., 2021). Similar to other metrics, Averaging these scores across prompts yields the final score.
@@ C.1 State-of-the-art Loss Functions
The proposed loss functions of three well-knwon training-free approach, Attant-and-Excite (Chefer et al., 2023), Divide-and-Bind (Li et al., 2023), and Predicated-Diffusion (Sueyoshi & Matsubara, 2024), are demonstrated in equations 8, 9, and 10, respectively. At each denoising step $t$, the latent embedding $z_{t}$ is modified in the direction opposite to the gradient of these loss functions with respect to $z_{t}$.
@@ D.2 Results of Free-Style Prompts
In the previous experiments, prompts followed a fixed format of “a {entity} and a {entity}”. To evaluate the generalizability of our method to more diverse linguistic structures, we also tested it on free-form captions from the COCO dataset (Lin et al., 2014). Table 12 presents the results for COCO captions containing at least three entities, utilizing Stable Diffusion 1.4 as the backbone model. The overlap-based loss functions demonstrate significant improvements over the baseline methods.
@@ D.3 Results of Other Backbones
In addition to SD v1.4, we assessed our approach on other recent backbones, SD v2-base, SD v2 (Rombach et al., 2022), and SD-XL (Podell et al., 2023), to validate its effectiveness across different model sizes. Tables 13 and 14 present the results on two-entity and three-entity prompts from our COCO-Comp dataset using Stable Diffusion v2-base. Moreover, tables 15 and 16 demonstrate the obtained results on two-entity prompts of the COCO-Comp dataset, using Stable Diffusion v2 and Stable Diffusion XL, respectively. Across all the backbones, the overlap-based methods outperform other baselines in all evaluation metrics. This suggests that our overlap-based approach can generalize well to different diffusion model architectures and sizes.
@@ Appendix E Additional Qualitative Results
Figures 7, 8, 9, 10, 11, and 12 present additional qualitative comparisons of images generated between our overlap-based methods and various baselines. More specifically, figures 7, 10, 11, and 12 contain two-object prompts from our COCO-Comp dataset using Stable Diffusion XL, Stable Diffusion v2, Stable Diffusion v1.4, Stable Diffusion v2-base, respectively. Moreover, figure 8 includes free-style and complex prompts from the COCO dataset. Finally, figure 9 demonstrates the generated images from prompts containing background entities such as ‘sea’, ‘beach’, ‘desert’, and ‘mountain’.
Figure 7: Qualitative comparison between our proposed overlap-based methods and the state-of-the-art baseline, Stable Diffusion XL, on diverse prompts from our COCO-Comp dataset.
Figure 8: Qualitative comparison between our proposed overlap-based method, CoM distance, and the state-of-the-art baseline, Stable Diffusion XL, on free-style and complex prompts from the COCO dataset.
Figure 9: Qualitative comparison between our proposed overlap-based method, CoM distance, and the state-of-the-art baseline, Stable Diffusion XL, on prompts containing background entities such as ‘sea’, ‘beach’, ‘desert’, and ‘mountain’.
@@ Appendix G Limitations
The poor capability of the CLIP text encoder (Radford et al., 2021) in understanding and representing the compositionality of complex texts (Yuksekgonul et al., 2023; Thrush et al., 2022) may have a significant impact on the performance of T2I models. Since most of these models rely on the CLIP text encoder as their backbone for processing textual inputs, any limitations or weaknesses in the encoder’s ability to capture and encode the compositional structure and semantics of the text can propagate through the entire system, leading to suboptimal results or failure modes in the generated images. Surprisingly, most of the related works have not paid sufficient attention to this potential root cause of the failures. Instead, they often focus on proposing solutions without thoroughly investigating the underlying reasons behind the problems. They tend to attribute the issues to other factors without considering the limitations of the CLIP text encoder itself. To address this, we plan to focus our future work on investigating and improving the text encoding capabilities of T2I models, specifically targeting their ability to handle compositional complexities in text prompts.